# Variable Rate Learned Wavelet Video Coding with Temporal Layer Adaptivity

### v1

@@ IV-B Comparison to State-of-the-Art Video Coders
DCVC-HEM [20] -8.44% -11.49% +6.66 % +22.88%
DCVC-DC [15] -15.06% -14.95% +1.63 % +21.00%
pMCTF-CA [7] +0.97% +81.45% +3.26% +91.38%
pMCTF-CA-L (ours) -17.49% -25.04% -8.73% -24.83%
VTM LD-P -91.36% -55.81% -10.68 % -31.89%
To evaluate state-of-the-art learned coders, we use DCVC-HEM [20] and DCVC-DC [15] with a GOP of 8 for a fair comparison to our MCTF approach. Because the GOP structure of MCTF is sensitive to the choice of GOP size, we allow a content adaptive GOP choice from [6] and obtain our model pMCTF-CA-L. In addition, we include other learned MCTF approaches in our experiments, namely MCTF-CA [6] and pMCTF-CA [7]. The latter has a parallelized context model in comparison to MCTF-CA, but both models have no variable rate support and were trained with a single temporal layer only. To consider a traditional hybrid video coder, we include VTM 17.2 [22]. We use VTM in Lowdelay P (LD-P) configuration with an intra period of 8 such that it performs unidirectional motion estimation and has the same I frame number as the learned coders.
The rate-distortion curves on the UVG data set can be seen in Fig. 4a. The curves demonstrate that our pMCTF-CA-L model performs especially well for high rates. Here, our approach is the only learning-based coder that outperforms VTM at bitrates above $\approx 0.11$ bpp. Because of the model’s capability of lossless compression, it has the capacity to perform specifically well for higher rates as well. In contrast, the models DCVC-HEM and DCVC-DC have better performance at low rates. For the UVG 4K data set, the rate-distortion curves in Fig. 4b show a similar behavior. Our pMCTF-CA-L model clearly outperforms the other MCTF models and also performs better than DCVC-HEM and DCVC-DC.
For a quantitative evaluation of the rate-distortion performance, we measure average Bjøntegaard Delta (BD) rate savings of different coders with MCTF-CA as an anchor. The results are shown in Table II. We find that our pMCTF-CA-L model outperforms the other learned coders in terms of PSNR as well as MS-SSIM on both the UVG and UVG 4K data set. In particular, we outperform DCVC-DC in terms of PSNR, which achieves BD rate savings of -15.06%, while our model gives savings of -17.49% on the UVG data set. On the UVG 4K data set, our model is about 10% better than DCVC-DC. The BD rate savings show that VTM outperforms the learning-based coders overall, whereby the performance gap is smaller for MS-SSIM.
IV-C Ablation Study
@@
We demonstrate the effectiveness of our approach by an ablation study on the UVG data set. We take the pMCTF model from [7] as baseline. Fig. 4c provides the rate-distortion curves of the baseline pMCTF model and different configurations of our proposed method with BD rate savings in comparison to the baseline. The brown curve ”pMCTF var. rate” achieves average BD rate savings of -10.3% by enabling variable rate support, fixing the weights of the optical flow network, and using a four-step context model for motion vector coding. The model ”pMCTF var. rate + two lvls” gives BD rate savings of -16.5% by adding training stage six (cf. Table I) with two temporal levels. With the temporal layer adaptive quality scaling, we get the model ”pMCTF var. rate + two lvls + $q_{\mathrm{scale}}^{(j)}$” (pMCTF-L) achieving BD rate savings of -19.4% and hence further increasing coding efficiency. The red curve corresponding to pMCTF-CA-L with layer adaptivity and a content adaptive GOP choice achieves average BD rate savings of -23.0% over the baseline pMCTF model.
IV-D Complexity
@@
TABLE III: Complexity comparison in terms of model size, kMAC/px, and decoding time. Runtime is measured on a NVIDIA RTX A6000 GPU on center crops of size $512\times 768$ from the UVG sequences.
Model Parameter size kMAC/px Decoding time
MCTF-CA [6] 90 MB 3554 1865.6s
pMCTF-CA [7] 182 MB 6593 7.7s
pMCTF-CA-L (ours) 144 MB 5357 8.7s
@@
In Table III, we assess the complexity of our method compared to other MCTF-based approaches. To this end, we measure average decoding times, complexity in terms of kilo multiply-accumulate operations per pixel (kMAC/px), and provide the model size. Regarding the model size, the pMCTF models are larger due to the four-step context fusion models. We decreased the model size of our pMCTF-L compared to pMCTF by reducing the number of channels within the context fusion model. This results in less kMAC/px for our pMCTF-L model compared to pMCTF as well. In terms of kMAC/px, our model is more complex than the MCTF model from [6], however, the average decoding time is significantly lower, which makes our pMCTF-L model most practical. Also, our model only requires a single checkpoint for multiple rate-distortion points.
V Conclusion
@@
In this paper, we introduced a training strategy that enables learned MCTF coders to consider two temporal layers during training. Thereby, we reduced the train-test gap and enabled temporal layer adpative quality scaling with flexible rate support. Overall, our proposed coder pMCTF-CA-L achieves BD rates savings of at least -17% compared to pMCTF-CA and also outperforms state-of-the-art learned coders like DCVC-DC. In future work, we will explore ways of considering information from multiple reference frames in the MCTF framework, e.g., by supporting bidirectional prediction.

### v3

@@ 4.2 Comparison to State-of-the-Art Video Coders
VTM RA -26.42% -28.6% -19.34% -29.9%
DCVC-HEM [20] 36.0% 158.2% 35.4% 195.5%
DCVC-FM [17] -43.7% -35.7% -36.1 % -40.5%
Cetin et al. [22] 15.1% 630% 40.6% 541.6%
MCTF-CA [6] 51.0 % 192.30% 21.5% 104.7%
pMCTF-CA [7] 56.9% 390.2% 31.3% 283.5%
pMCTF-CA-L (ours) -8.0% 12.0% -20.0% -33.0%
Fig. 5: Ablation study on the UVG data set.
4.3 Ablation Study
@@
We demonstrate the effectiveness of our approach by an ablation study on the UVG data set at a GOP size of 8. We take the pMCTF model from [7] as baseline which was designed for a maximum GOP size of 8. Fig. 5 provides the rate-distortion curves of the baseline pMCTF model and different configurations of our method. The brown curve ”pMCTF var.” achieves average BD rate savings of -10.3% by enabling variable rate support, fixing the weights of the optical flow network, and using a four-step context model for motion vector coding. The model ”pMCTF var. + two lvls” gives BD rate savings of -16.5% by training on two temporal levels. With the temporal layer adaptive quality scaling, we get the model ”pMCTF var. + two lvls + $q_{\mathrm{scale}}^{(j)}$” achieving BD rate savings of -19.4%. When further increasing the training sequence length to 8 and enabling a latent prior for motion vector coding, we obtain our model pMCTF-L with BD rate savings of -26.6%. The red curve corresponding to pMCTF-CA-L with layer adaptivity and a content adaptive GOP choice achieves average BD rate savings of -34.6% over the baseline pMCTF model.
4.4 Complexity
@@
Table 3: Complexity comparison in terms of kMAC/px and decoding time. Runtime is measured on a NVIDIA RTX A6000 GPU on center crops of size $512\times 768$ from the UVG sequences.
Model kMAC/px Decoding time
MCTF-CA [6] 3554 1865.6s
pMCTF-CA [7] 6593 7.7s
pMCTF-CA-L (ours) 5401 8.7s
@@
In Table 3, we assess the complexity of our method compared to other MCTF-based approaches. To this end, we measure average decoding times and complexity in terms of kilo multiply-accumulate operations per pixel (kMAC/px). The pMCTF models require more MACs due to the four-step context fusion models. By reducing the number of channels within this context model, our pMCTF-L model needs less MACs than pMCTF. Our model is more complex than the MCTF model from [6], however, the average decoding time is significantly lower, which makes our pMCTF-L model most practical.
5 Conclusion
@@
In this paper, we introduced variable rate support for learned MCTF and temporal layer adaptive quality scaling. By presenting a training strategy that supports multiple temporal layers we closed the train-test gap for learned MCTF coders. Overall, our proposed coder pMCTF-CA-L achieves BD rates savings of at least -32% compared to pMCTF-CA for a GOP size of 16. This makes our coder competitive to VTM LD-P and advances learned wavelet video coders as an alternative learned video coding scheme with a known structure. In future work, we will explore ways of considering information from multiple reference frames in the MCTF framework.